# Numerical linked cluster expansions for two-dimensional spin models with continuous disorder distributions

### v1

@@ Abstract
We show that numerical linked cluster expansions (NLCEs) based on sufficiently large building blocks can be used to obtain accurate low-temperature results for the thermodynamic properties of spin lattice models with continuous disorder distributions. Our main finding is that such results can be obtained using direct disorder averages on the NLCE clusters before computing their weights. We provide a proof of concept using three different NLCEs based on L, square, and rectangle building blocks. We consider both classical (Ising) and quantum (Heisenberg) spin-$\frac{1}{2}$ models and show that convergence can be achieved down to temperatures that are up to two orders of magnitude lower than the relevant energy scale in the model. Additionally, we provide evidence that in one dimension one can compute observables such as the energy down to their ground state values.
@@ I Introduction
Disorder, resulting from lattice impurities, distortions, or vacancies, can alter the properties of materials in a drastic fashion. For example, noninteracting electrons in the presence of disorder can exhibit Anderson localization Anderson . In spin models, the focus of our work, quenched disorder – affecting the spin exchange interactions – can lead to frustration and spin glasses. Frustration can preclude spin ordering as the temperature is decreased and, below a critical temperature whose value depends on the specific model, a spin glass may form Edwards_1975 ; Binder_spinglass ; Weissman_spinglass .
@@
In this work, we show that numerical linked cluster expansions (NLCEs) can be used to obtain accurate low-temperature results for the thermodynamic properties of classical and quantum spin models with continuous disorder distributions. Previous studies have already shown that NLCEs can be used to obtain accurate results for bimodal Tang_dynamicdis ; Tang_BiDisorder and multimodal Mulanix_mltidis ; park_khatami_21 disorder distributions, and that increasing the number of modes in properly selected multimodal disorder distributions can be used to approximate the results for continuous disorder distributions Mulanix_mltidis ; park_khatami_21 . Our goal here is to show that NLCEs based on large building blocks, such as Ls, squares, and rectangles can be used to carry out direct samplings of continuous disorder distributions to produce accurate results for thermodynamic properties at low temperatures.
@@
The same applies, in principle, to continuous disorder distributions. However, for continuous distributions, it is impossible to carry out the exact disorder averages except for small clusters. Averages over finite numbers of disorder realizations carry statistical errors that result in a divergence of the NLCEs for the commonly used bond and site expansions introduced in Refs. rigol_bryant_06 ; rigol_bryant_07a . Divergences occur because computing the weights of large clusters in such expansions involves subtracting weights of exponentially many smaller subclusters, whose statistical errors add up. An alternative way to proceed is to carry out subtractions directly for any given disorder realization on any given cluster (there are no statistical errors in that case) and then average over disorder realizations for that cluster devakul_15 ; Hazzard_dis . This is computationally very demanding and has yet to be successfully implemented in the context of thermodynamic properties of quantum models with continuous disorder distributions.
@@
The presentation is organized as follows. The spin-$\frac{1}{2}$ Ising and Heisenberg Hamiltonians studied in this work are introduced in Sec. II. A brief review of NLCEs, the resummation algorithms used, and the observables that we calculate is provided in Sec. III. To build up to our work in the two-dimensional (2D) square lattice, in Sec. IV we use NLCEs to find a closed form expression for the thermodynamic properties of the one-dimensional (1D) Ising model with an arbitrary disorder distribution, as well as to numerically study the 1D Heisenberg model in the presence of a uniform disorder distribution. The specific expansions used here in 2D – the restricted L, the square, and the rectangle expansions – are discussed in Sec. V. In Sec. VI, we compare results obtained using those three expansions against site-expansion results reported in Ref. Tang_BiDisorder for the Ising model with a bimodal disorder distribution. The results for the Ising and Heisenberg models with uniform disorder distributions are reported in Sec. VII. We conclude with a summary and discussion of our results in Sec. VIII.
@@ II Model Hamiltonians
where $\hat{S}^{z}_{\bf i}$ is the $z$-component of the spin-$\frac{1}{2}$ operator at site ${\bf i}$, and $\langle{\bf i},{\bf j}\rangle$ denotes pairs of nearest neighbors sites. Note that the interaction strength $J_{\bf ij}$ depends on the pair of sites $\langle{\bf i},{\bf j}\rangle$. We draw $J_{\bf ij}$ from different discrete and continuous disorder distributions, as specified later.
@@ III A short summary of NLCEs
NLCEs allow one to calculate finite-temperature properties of extensive observables for translationally invariant lattice models in the thermodynamic limit. For an extensive observable $\mathcal{O}$, its corresponding intensive counterpart per lattice site $O\equiv\mathcal{O}/N$ can be computed using the following sum over all the connected clusters that can be embedded on the lattice
@@
In order to evaluate Eq. (3), the sum is truncated and one only includes the clusters that can be solved numerically. Convergence at any given temperature $T$ is achieved when the results of successive orders, which are determined by the largest clusters considered, agree with each other. For unordered phases, the NLCE results have been shown to approach the thermodynamic limit results exponentially fast in the NLCE order iyer_15 .
Because of the lack of translational invariance in models with disorder, one may think that completely independent NLCE calculations need to be carried out for each disorder realization, so that weights can be properly subtracted. However, as noted in Ref. Tang_BiDisorder in the context of bimodal disorder distributions, averaging over all possible disorder realizations restores translational invariance. This, together with the linear character of the NLCE Eqs. (3) and (4), allows one to use the exact same expansion as for translationally invariant models. We use that approach here. Namely, we use Eqs. (3) and (4) after replacing the expectation values of the observables in each cluster $c$ by their disorder averages $\overline{\mathcal{O}(c)}$. For discrete disorder distributions, such as bimodal disorder, the averages can be computed exactly Tang_BiDisorder . For continuous disorder distributions, we set a maximum value of the standard deviation for all clusters of any given size, and explore the effect that changing such a maximum (which may depend on the cluster size) has on the NLCE results.
@@
$\hat{\rho}=\frac{1}{Z}\exp\left(-\frac{\hat{H}}{k_{B}T}\right),\ \text{with}\ Z=\text{Tr}\left[\exp\left(-\frac{\hat{H}}{k_{B}T}\right)\right]\,.$ (5)
$\hat{H}$ is the model Hamiltonian, $k_{B}$ is the Boltzmann constant (we set $k_{B}=1$), and $T$ is the temperature (which has units of energy in our convention). We compute three thermodynamic quantities the energy $E$, the entropy $S$, and the specific heat $C_{v}$, all per site.
To gauge the convergence of the direct sums in Eq. (3), we calculate the normalized difference for each order $l$ with respect to the highest order $l_{\text{max}}$ accessible to us
@@
In order to obtain results at temperatures lower than those at which the direct sums converge, we use resummation techniques. Specifically, we use two resummation techniques rigol_bryant_07a :
@@
where $k$ denotes the number of Wynn resummation “cycles”. Only even entries $\epsilon_{l}^{(2k^{\prime})}$ (with $k^{\prime}$ an integer) are expected to converge to the thermodynamic limit result. We note that the new sequence generated after two cycles has two fewer terms. The estimate for an observable after $2k^{\prime}$ cycles is given by
@@
(ii) Euler algorithm, which can accelerate the convergence of alternating series. In this algorithm, see Ref. Tang_2013 , the only free parameter is the number of terms “$k$” for which the direct sum is carried out before the Euler transformation is used. In what follows whenever we report the results of the Euler algorithm, $\text{Euler}_{k}(O)$, we specify the value of $k$ used.
@@ IV Ising and Heisenberg models in 1D
In this section, we study the thermodynamic properties of the Ising and Heisenberg models with continuous disorder distributions in 1D. We note that, in 1D, pairs of nearest neighbors sites $\langle{\bf i},{\bf j}\rangle\equiv i,i+1$, i.e., we can parameterize the bonds $J_{\bf ij}$ with one index and write $J_{i}$.
@@ IV.1.1 Exact Solution
For an open chain with $N$ sites, using the traditional transfer matrix method it is straightforward to calculate the partition function $Z_{N}=2\prod_{i=1}^{N-1}2\cosh(\beta J_{i})$, where $\beta=1/T$ (we set $k_{B}=1$). One can therefore calculate the intensive quantity $\ln(Z_{N})/N$, and average over the PDF in the limit $N\rightarrow\infty$
@@
$+\lim_{N\rightarrow\infty}\frac{ln(2)}{N}.$
@@
IV.1.2 NLCE solution
For a single site we only have one configuration, with 2 possible states, so the partition function is trivially $\ln(Z_{1})=\ln 2$. For 2 sites, the partition function is $Z_{2}=2(e^{\beta J}+e^{-\beta J})$ where $J$ is a random coupling constant chosen from the PDF $P(J)$. We then average over all possible $J$’s to get the average $\ln(Z_{2})$
@@ IV.1.2 NLCE solution
For the open chain with 3 sites, there are two different coupling constants, $J_{1}$ and $J_{2}$, drawn from identical independent PDFs $P(J)$. The partition function follows
@@
which is the exact solution, see Eq. (10). Hence, like for the clean Ising model iyer_15 , the second order NLCE provides an exact solution for the 1D Ising model with an arbitrary disorder distribution.
@@ IV.2 Heisenberg model
In contrast to the translationally invariant case, the (quantum) 1D Heisenberg model with a continuous disorder distribution is not exactly solvable Ma_79 ; Dasgupta_Ma_80 . We study this model numerically using NLCEs, and focus on the zero-mean uniform disorder distribution, with PDF
@@
In 1D models with only nearest neighbor couplings, there is one cluster at each order $l$ of the NLCE; an open chain with $l$ sites. When calculating the $l_{\text{th}}$ order NLCE result for an observable, Eq. (3) simplifies to $O_{l}=\overline{\mathcal{O}(l)}-\overline{\mathcal{O}(l-1)}$. So, one only needs to calculate observables for two consecutively cluster sizes in order to get the NLCE result at any given order. In Fig. 1, we plot NLCE results for the energy $E$, the entropy $S$, and the specific heat $C_{v}$ vs $T$ for the 1D Heisenberg model with a uniform disorder distribution. We set $J=1$ to be our energy scale and report results for $l=13$ and $l=14$. Those results were obtained carrying out averages over 45, 17.5, and 4 million disorder realizations for the chains with 12, 13, and 14 sites, respectively. We also show the corresponding results for the clean model with $J=1$.
@@
Figure 1: Thermodynamic properties of the 1D Heisenberg model with a uniform disorder distribution, and for the clean Heisenberg model. We plot the NLCE results for (a) energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ vs $T$ obtained at orders $l=13$ and 14 ($l=17$ and 18) for the model with disorder (clean model), and Wynn’s and Euler’s resummation results for the clean model. The horizontal dotted line in (a) shows the exact result for the ground state energy of the clean model, $E_{G}=1/4-\ln 2\approx-0.443$ Giamarchi_03 .
@@
For the energy in the presence of disorder [Fig. 1(a)], we find that the results for $l=13$ and $l=14$ agree with each other down to $T=10^{-3}$, at which $E$ has become temperature independent and we essentially obtain the ground-state energy. This is to be contrasted to the results for the clean model, for which the results for $l=17$ and $l=18$ agree with each other down only to $T\approx 0.2$. Disorder, which reduces correlations, extends the NLCE convergence to lower temperatures for all the observables considered here. Resummations for the clean model do allow one to reproduce the exact ground state energy (shown as a horizontal dotted line). For the entropy in the presence of disorder [Fig. 1(b)], the results for $l=13$ and $l=14$ agree with each other down to $T\approx 10^{-2}$, in comparison to $T\approx 0.2$ for the clean model. The contrast between the results in the presence and absence of disorder make apparent that the relatively high value of the entropy at $T\approx 10^{-2}$ in the former is a consequence of frustration introduced by the random couplings. In the clean model, the resummation results indicate that the entropy at that temperature is vanishingly small. Like for the entropy, for the specific heat in the presence of disorder [Fig. 1(c)], the results for $l=13$ and $l=14$ agree with each other down to $T\approx 10^{-2}$. NLCEs show that there is a well-resolved peak in the specific heat with a maximum value $C^{\text{max}}_{v}\approx 0.23$ at $T_{m}\approx 0.25$. Disorder can be seen to reduce the height, and move towards lower temperatures, the peak appearing in the clean model, which has a maximum $C^{\text{max}}_{v}\approx 0.37$ at $T_{m}\approx 0.45$.
In Fig. 2, we compare the differences
$D_{13}(O)=|O_{13}-O_{14}|,$ (17)
to the statistical errors
$\sigma_{l}(O)=\sqrt{\text{Var}[\mathcal{O}(l)]+\text{Var}[\mathcal{O}(l-1)]},$ (18)
for $l=13$ and 14. We report results for the energy and the specific heat (the results for the entropy, not shown, are qualitatively similar to those for the specific heat). We do not normalize $D_{13}(O)$ nor $\sigma_{l}(O)$ because the energy vanishes at $T=\infty$, and the specific heat vanishes at $T=0$ and $T=\infty$.
@@
Figure 2: Errors in the NLCE calculations of the thermodynamic properties of the 1D Heisenberg model with a uniform disorder distribution. Differences $D_{13}(O)$, see Eq. (17), and the statistical errors $\sigma_{l}(O)$, see Eq. (18), for $l=13$ and 14 for the (a) energy $E$, and (b) specific heat $C_{v}$.
@@
For the energy in Fig. 2(a), one can see that $D_{13}(E)$ is of the order of (or smaller than) the statistical errors at all temperatures. This shows that the statistical errors are the dominant errors in our NLCE calculations of the energy at all temperatures. For the specific heat, the results in Figs. 2(b) show that the statistical errors are dominant only at temperatures $T\gtrsim 0.03$. At very low temperatures $T\lesssim 0.03$, $D_{13}(C_{v})$ becomes much larger than the statistical errors, which shows that lack of convergence dominates the error in the NLCE calculations in that regime (as it does in the clean case).

### v2

# Numerical linked-cluster expansions for two-dimensional spin models with continuous disorder distributions

@@ Abstract
We show that numerical linked cluster expansions (NLCEs) based on sufficiently large building blocks allow one to obtain accurate low-temperature results for the thermodynamic properties of spin lattice models with continuous disorder distributions. Specifically, we show that such results can be obtained computing the disorder averages in the NLCE clusters before calculating their weights. We provide a proof of concept using three different NLCEs based on L, square, and rectangle building blocks. We consider both classical (Ising) and quantum (Heisenberg) spin-$\frac{1}{2}$ models and show that convergence can be achieved down to temperatures that are up to two orders of magnitude lower than the relevant energy scale in the model. Additionally, we provide evidence that in one dimension one can obtain accurate results for observables such as the energy down to their ground-state values.
@@ I Introduction
Disorder, resulting from lattice impurities, distortions, or vacancies, can alter the properties of materials in a drastic fashion. For example, noninteracting electrons in the presence of disorder can exhibit Anderson localization Anderson . In spin models, the focus of our work, quenched disorder—affecting the spin exchange interactions—can lead to frustration and spin glasses. Frustration can preclude spin ordering as the temperature is decreased and, below a critical temperature whose value depends on the specific model, a spin glass may form Edwards_1975 ; Binder_spinglass ; Weissman_spinglass .
@@
In this work, we show that numerical linked-cluster expansions (NLCEs) can be used to obtain accurate low-temperature results for the thermodynamic properties of classical and quantum spin models with continuous disorder distributions. Previous studies have already shown that NLCEs can be used to obtain accurate results for bimodal Tang_dynamicdis ; Tang_BiDisorder and multimodal Mulanix_mltidis ; park_khatami_21 disorder distributions and that increasing the number of modes in properly selected multimodal disorder distributions can be used to approximate the results for continuous disorder distributions Mulanix_mltidis ; park_khatami_21 . Our goal here is to show that NLCEs based on large building blocks, such as Ls, squares, and rectangles can be used to carry out direct samplings of continuous disorder distributions to produce accurate results for thermodynamic properties at low temperatures.
@@
The same applies, in principle, to continuous disorder distributions. However, for continuous distributions, it is impossible to carry out the exact disorder averages except for small clusters. Averages over finite numbers of disorder realizations carry statistical errors that result in a divergence of the NLCEs for the commonly used bond and site expansions introduced in Refs. rigol_bryant_06 ; rigol_bryant_07a . Divergences occur because computing the weights of large clusters in such expansions involves subtracting weights of exponentially many smaller subclusters, whose statistical errors add up. An alternative way to proceed is to carry out subtractions directly for any given disorder realization on any given cluster (there are no statistical errors in that case) and then average over disorder realizations for that cluster devakul_15 ; Hazzard_dis . This is computationally very demanding and has yet to be successfully implemented in the context of thermodynamic properties of quantum models with continuous disorder distributions at finite temperature (see Ref. Pardini_19 for ground-state calculations).
@@
The presentation is organized as follows. The spin-$\frac{1}{2}$ Ising and Heisenberg Hamiltonians studied in this work are introduced in Sec. II. A brief review of NLCEs, the resummation algorithms used, and the observables that we calculate is provided in Sec. III. To build up to our work in the two-dimensional (2D) square lattice, in Sec. IV we use the linked cluster theorem to find a closed form expression for the thermodynamic properties of the 1D Ising model with an arbitrary disorder distribution, as well as NLCEs to numerically study the 1D Heisenberg model in the presence of a uniform disorder distribution. The specific expansions used here in 2D–the restricted L, the square, and the rectangle expansions–are discussed in Sec. V. In Sec. VI, we compare results obtained using those three expansions against site-expansion results reported in Ref. Tang_BiDisorder for the Ising and Heisenberg models with a bimodal disorder distribution. The results for the Ising and Heisenberg models with uniform disorder distributions are reported in Sec. VII. We conclude with a summary and discussion of our results in Sec. VIII.
@@ II Model Hamiltonians
where $\hat{S}^{z}_{\bf i}$ is the $z$ component of the spin-$\frac{1}{2}$ operator at site ${\bf i}$ and $\langle{\bf i},{\bf j}\rangle$ denotes pairs of nearest-neighbors sites. Note that the interaction strength $J_{\bf ij}$ depends on the pair of sites $\langle{\bf i},{\bf j}\rangle$. We draw $J_{\bf ij}$ from different discrete and continuous disorder distributions, as specified later.
@@ III A short summary of NLCEs
NLCEs allow one to calculate finite-temperature properties of extensive observables for translationally invariant lattice models in the thermodynamic limit. For an extensive observable $\mathcal{O}$, its corresponding intensive counterpart per lattice site $O\equiv\mathcal{O}/N$ can be computed using the linked cluster theorem, namely using the following sum over all the connected clusters that can be embedded on the lattice:
@@
In NLCEs, one truncates the sum in Eq. (3) to include only the clusters that can be solved exactly numerically. Convergence at any given temperature $T$ is achieved when the results of successive orders, labelled by the largest clusters considered, agree with each other. For unordered phases, the NLCE results have been shown to approach the thermodynamic limit results exponentially fast in the NLCE order iyer_15 .
Because of the lack of translational invariance in models with disorder, one may think that completely independent NLCE calculations need to be carried out for each disorder realization, so that weights can be properly subtracted. However, as noted in Ref. Tang_BiDisorder in the context of bimodal disorder distributions, averaging over all possible disorder realizations restores translational invariance. This, together with the linear character of the NLCE Eqs. (3) and (4), allows one to use the exact same expansion as for translationally invariant models. We use that approach here. Namely, we use Eqs. (3) and (4) after replacing the expectation values of the observables in each cluster $c$ by their disorder averages $\overline{\mathcal{O}(c)}$. For discrete disorder distributions, such as bimodal disorder, the averages can be computed exactly Tang_BiDisorder . For continuous disorder distributions, we set a maximum value of the normalized standard deviation for all clusters of any given size and explore the effect that changing such a maximum (which in general depends on the cluster size) has on the NLCE results.
@@
$\hat{\rho}=\frac{1}{Z}\exp\left(-\frac{\hat{H}}{k_{B}T}\right),\ \text{with}\ Z=\text{Tr}\left[\exp\left(-\frac{\hat{H}}{k_{B}T}\right)\right]\,,$ (5)
where $\hat{H}$ is the model Hamiltonian, $k_{B}$ is the Boltzmann constant (we set $k_{B}=1$), and $T$ is the temperature (which has units of energy in our convention). We compute three thermodynamic quantities, the energy $E$, the entropy $S$, and the specific heat $C_{v}$, all per site.
To gauge the convergence of the direct sums in Eq. (3), we calculate the normalized difference for each order $l$ with respect to the highest-order $l_{\text{max}}$ accessible to us
@@
In order to obtain results at temperatures lower than those at which the direct sums converge, we use resummation techniques. Specifically, we use the following two resummation techniques rigol_bryant_07a :
@@
where $k$ denotes the number of Wynn resummation “cycles.” Only even entries $\epsilon_{l}^{(2k^{\prime})}$ (with $k^{\prime}$ an integer) are expected to converge to the thermodynamic limit result. We note that the new sequence generated after two cycles has two fewer terms. The estimate for an observable after $2k^{\prime}$ cycles is given by
@@
(ii) The Euler algorithm, which can accelerate the convergence of alternating series. In this algorithm, see Ref. Tang_2013 , the only free parameter is the number of terms “$k$” for which the direct sum is carried out before the Euler transformation is used. In what follows whenever we report the results of the Euler algorithm, $\text{Euler}_{k}(O)$, we specify the value of $k$ used.
@@ IV Ising and Heisenberg models in 1D
In this section, we study the thermodynamic properties of the Ising and Heisenberg models with continuous disorder distributions in 1D. We note that, in 1D, pairs of nearest-neighbors sites $\langle{\bf i},{\bf j}\rangle\equiv i,i+1$, i.e., we can parametrize the bonds $J_{\bf ij}$ with one index and write $J_{i}$.
@@ IV.1.1 Exact solution
For an open chain with $N$ sites, using the traditional transfer matrix method it is straightforward to calculate the partition function $Z_{N}=2\prod_{i=1}^{N-1}2\cosh(\beta J_{i})$, where $\beta=1/T$ (we set $k_{B}=1$). One can therefore calculate the intensive quantity $\ln(Z_{N})/N$ and average over the PDF in the limit $N\rightarrow\infty$,
@@
$+\lim_{N\rightarrow\infty}\frac{\ln 2}{N}.$
@@
IV.1.2 Linked cluster theorem solution
For a single site we only have one configuration, with two possible states, so the partition function is trivially $\ln(Z_{1})=\ln 2$. For two sites, the partition function is $Z_{2}=2(e^{\beta J}+e^{-\beta J})$, where $J$ is a random coupling constant chosen from the PDF $P(J)$. We then average over all possible $J$’s to get the average $\ln(Z_{2})$
@@ IV.1.2 Linked cluster theorem solution
For the open chain with three sites, there are two different coupling constants, $J_{1}$ and $J_{2}$, drawn from identical independent PDFs $P(J)$. The partition function follows
@@
which is the exact solution, see Eq. (10). Hence, like for the clean Ising model iyer_15 , the linked cluster theorem allows one to find the exact solution for the 1D Ising model with an arbitrary disorder distribution.
@@ IV.2 Heisenberg model
In contrast to the translationally invariant case, the (quantum) 1D Heisenberg model with a continuous disorder distribution is not exactly solvable Ma_79 ; Dasgupta_Ma_80 . We study this model numerically using NLCEs and focus on the zero-mean uniform disorder distribution, with PDF
@@
In 1D models with only nearest-neighbor couplings, there is one cluster at each order $l$ of the NLCE; an open chain with $l$ sites. When calculating the $l\text{th}$ order NLCE result for an observable, Eq. (3) simplifies to $O_{l}=\overline{\mathcal{O}(l)}-\overline{\mathcal{O}(l-1)}$. So one only needs to calculate observables for two consecutively cluster sizes in order to get the NLCE result at any given order. In Fig. 1, we plot NLCE results for the energy $E$, the entropy $S$, and the specific heat $C_{v}$ vs $T$ for the 1D Heisenberg model with a uniform disorder distribution. We set $J=1$ to be our energy scale and report results for $l=14$ and $l=15$. Those results were obtained carrying out averages over 17.5, 4, and 0.6 million disorder realizations for the chains with 13, 14, and 15 sites, respectively. We also show the corresponding results for the clean model with $J=1$.
@@
Figure 1: Thermodynamic properties of the 1D Heisenberg model with a uniform disorder distribution, and for the clean Heisenberg model. We plot the NLCE results for (a) energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ vs $T$ obtained at orders $l=14$ and 15 ($l=17$ and 18) for the model with disorder (clean model), and Wynn’s and Euler’s resummation results for the clean model. The horizontal dotted line in (a) shows the exact result for the ground-state energy of the clean model, $E_{G}=1/4-\ln 2\approx-0.443$ Giamarchi_03 .
@@
For the energy in the presence of disorder [Fig. 1(a)], we find that the results for $l=14$ and $l=15$ agree with each other down to $T=10^{-3}$, at which $E$ has become temperature independent and we essentially obtain the ground-state energy. This is to be contrasted to the results for the clean model, for which the results for $l=17$ and $l=18$ agree with each other down only to $T\approx 0.2$. Disorder, which reduces correlations, extends the NLCE convergence to lower temperatures for all the observables considered here. Resummations for the clean model do allow one to reproduce the exact ground-state energy (shown as a horizontal dotted line). For the entropy in the presence of disorder [Fig. 1(b)], the results for $l=14$ and $l=15$ agree with each other down to $T\approx 10^{-2}$, in comparison to $T\approx 0.2$ for the clean model. The contrast between the results in the presence and absence of disorder make apparent that the relatively high value of the entropy at $T\approx 10^{-2}$ in the former is a consequence of frustration introduced by the random couplings. In the clean model, the resummation results indicate that the entropy at that temperature is vanishingly small. Like for the entropy, for the specific heat in the presence of disorder [Fig. 1(c)], the results for $l=14$ and $l=15$ agree with each other down to $T\approx 10^{-2}$. NLCEs show that there is a well-resolved peak in the specific heat with a maximum value $C^{\text{max}}_{v}\approx 0.23$ at $T_{m}\approx 0.25$. On the other hand, the peak appearing in the clean model has a maximum $C^{\text{max}}_{v}\approx 0.37$ at $T_{m}\approx 0.45$. Disorder reduces the height of that peak and moves it towards lower temperatures.
Next, we compare convergence errors [see Eq. (6)] to normalized standard deviations
$\delta_{c}(\mathcal{O})=\frac{\sigma_{c}({\mathcal{O}})}{{\overline{\mathcal{O}(c)}}}\,,$ (17)
where
$\sigma_{c}({\mathcal{O}})=\sqrt{\overline{\mathcal{O}^{2}(c)}-\overline{\mathcal{O}(c)}^{2}},$ (18)
with ${\overline{\mathcal{O}(c)}}$ being the disorder average of the observable in cluster $c$, to gain insights on the temperatures at which lack of convergence due to the size of the clusters dominates over statistical errors, and vice versa.
@@
Figure 2: Errors in the NLCE calculations of thermodynamic properties of the 1D Heisenberg model with a uniform disorder distribution. Differences $\Delta_{14}(O)$, see Eq. (6), and the normalized standard deviations $\delta_{c}(\mathcal{O})$, see Eq. (17), for the chain clusters $c=14$ and 15 with 14 and 15 sites, respectively. (a) Energy $E$ and (b) specific heat $C_{v}$.
@@
In Fig. 2, we plot $\Delta_{14}$ and $\delta_{c}$ for the chain clusters with 14 and 15 sites (the two largest ones considered), both for the energy and the specific heat (the results for the entropy, not shown, are qualitatively similar to those for the specific heat). For the energy [Fig. 2(a)], $\Delta_{14}(E)$ is of the order of (slightly larger than) the normalized standard deviations for the two clusters at the temperatures shown. This makes apparent that the statistical errors are the main errors in the NLCE calculations of the energy at those temperatures. For the specific heat, the results in Figs. 2(b) show that the statistical errors are the main errors only at temperatures $T\gtrsim 0.01$ [at the temperatures at which the NLCE results for orders 14 and 15 are indistinguishable from each other in Fig. 1(c)]. At very low temperatures $T\lesssim 0.01$, $\Delta_{14}(C_{v})$ becomes much larger than the statistical errors, which shows that lack of convergence due to the size of the clusters dominates the error in the NLCE calculations in that regime (as it does in the clean case).